# Fair Wasserstein Coresets

### v3

@@ 4.2 Updating Feature Vectors (Line 5)
$\min_{\widehat{X}_{i}\in{\cal X}:i\in[m]}\sum_{i\in[n]}\sum_{j\in[m]}c({Z}_{i},\widehat{Z}_{j})P_{ij}$ (12)
for $P=P_{k}^{\star}$, in which each component of $P$ is nonnegative and $\widehat{Z}_{j}=(\widehat{d}_{j},\widehat{X}_{j},\widehat{y}_{j})$, for the known fixed $\widehat{d}_{i}$ and $\widehat{y}_{i}$. Furthermore, the matrix $P$ is sparse, containing at most $n$ non-zeros (as when updating $P_{k}^{\star}$ for problem (8), see Appendix A). Moreover, problem (12) can be separated into the following $m$ subproblems,
$\min_{\widehat{X}_{i}\in{\cal X}}\sum_{i\in[n]}c(Z_{i},\widehat{Z}_{j})P_{ij}\ \text{, \ for }j\in[m]\ .$ (13)
@@ 6 An Alternative View: Generalized clustering algorithm
The minimizer $P^{\star}$ of (17) can be written in closed form. For each $i\in[n]$, let $C_{ij_{i}^{\star}}$ denote a smallest component on the $i$-th row of $C$. Then the components of a minimizer $P^{\star}$ can be written as $P_{ij}^{\star}=\frac{1}{n}\cdot\mathbf{I}(j=j_{i}^{\star})$ (where $\mathbf{I}$ is the indicator function). Hence, without fairness constraints, FWC corresponds to Lloyd’s algorithm for clustering. Specifically, Lloyd’s algorithm iteratively computes the centroid for each subset in the partition and subsequently re-partitions the input based on the closeness to these centroids (Lloyd, 1982); these are the same operations FWC does in optimizing the surrogate function and solving problem (17). Thus, when $c(x,y)$ is correspondingly defined as $\|x-y\|_{1}$ or $\|x-y\|_{2}^{2}$, FWC corresponds to Lloyd’s algorithm applied to k-medians or k-means problems, except the centroids have fixed values for $\widehat{D}$ and $\widehat{Y}$ (see Section 3).
Comparison with k-means and k-medoids FWC and Lloyds’ algorithm for k-means or k-median share similar per-iteration complexity, with the main difference in complexity due to solving problem (8). We solve this problem efficiently by utilizing a variant of the FairWASP approach by Xiong et al. (2024) (Section 4.1), hence avoiding the usual complexity in solving optimal transport problems. As shown in Section 5, the leading term in the runtime complexity is $\mathcal{O}(nm)$, which comes from calculating and storing the cost matrix $C$. This level of complexity is the same as those in k-means and k-medoids. In addition, from our experiments we also see that the per-iteration complexity of FWC is roughly linear with the original dataset size $n$ (see the runtime experiment in Section 7 and Appendix C).
@@ 7 Experiments
We evaluate the performance of FWC  on 4 datasets widely used in the fairness literature (Fabris et al., 2022): (i) Adult (Becker & Kohavi, 1996), (ii) German Credit (Hofmann, 1994), (iii) Communities and Crime (Redmond, 2009) and (iv) Drug (Fehrman et al., 2017). For each dataset, we consider 3 different coreset sizes $m=5\%,10\%,20\%$ (apart from the Adult dataset, in which we select $m$ equal to 0.5%, 1% and 2% due to the large dataset size). We compare our approach with: (a) Fairlets and IndFair, two fair clustering approaches by Backurs et al. (2019) and Chhaya et al. (2022), (b) K-Median Coresets, a coreset approach by Bachem et al. (2018), (c) k-means (Lloyd, 1982) and k-medoids (Maranzana, 1963; Park & Jun, 2009), two classic clustering approaches and (d) Uniform Subsampling of the original dataset. For FWC, we consider three different values of the fairness violation hyper-parameters $\epsilon$ for the optimization problem in (5). We compute the fairness-utility tradeoff by first training a 2-layer multilayer perceptron (MLP) classifier with ReLu activations on the coresets created by each approach and then evaluating the classifier demographic disparity (fairness) and AUC (utility). Figure 1 shows the model with the best fairness-utility tradeoff across the three coreset sizes $m$, for each approach. FWC  obtains equal or better fairness-utility tradeoffs (smaller disparity at the same level of utility, higher utility with the same disparity, or both) across all datasets, and performance remains competitive even when using a fairness pre-processing approach (Kamiran & Calders, 2012). Appendix C.2 includes more experiments and details, which highlight that: (a) FWC  consistently achieves coresets that are closer in distribution to the original dataset with respect to the other methods and, although not natively minimizing clustering cost, also provide competitive performance for smaller datasets (Tables 3 and 4); (b) when added to the training data using the data augmentation schema proposed by Sharma et al. (2020, Section 2.1) FWC generally either increase the performance or reduce the demographic disparity in the downstream learning process.
@@
Figure 1: Top left: FWC runtime when changing the original dataset size $n$. Others: Fairness-utility tradeoff on real datasets for a downstream MLP classifier, selecting the model with the best fairness-utility tradeoff across three different coreset sizes $m$, with averages taken over 10 runs. FWC  consistently achieves a comparable/better tradeoff, even when adjusting the other coresets with a fairness pre-processing technique (Kamiran & Calders, 2012). See text and Appendix C.2 for more details.
Using FWC to improve fairness for LLM.
@@
Wang et al. (2023) evaluate GPT-3.5 and GPT-4 for fairness on predictive tasks for the UCI Adult dataset in a zero and few shot setting. We use a similar evaluation setup and use FWC in the few shot setting as examples and evaluate the results for the gender protected attribute. Specifically, we transform the tabular data into language descriptions, and ask GPT-3.5 Turbo and GPT-4 to perform classification tasks on it. We select $200$ samples to construct the test set and use a set of 16 samples found using FWC as examples. Further details on this experiment are provided in the appendix. The results are shown in Table 1. Examples provided by  FWC help reduce demographic disparity more than providing balanced few shot examples, while losing on predictive accuracy (note that the drop in accuracy is similar to the drop observed in Wang et al. (2023) and is representative of the fairness-utility trade-off). Owing to the token limitation of LLM’s, these representative coresets can evidently be valuable to provide a small set of samples that can help mitigate bias.
@@
Table 1: Using the same setup as in Wang et al. (2023), we use GPT-3.5 Turbo and GPT-4 LLM’s for fairness evaluations, with a test set of 200 samples with 0.5 base parity (${b_{p}}=0.5$). Few Shot - FWC is used to provide sixteen examples with weights to the model as examples. Accuracy and demographic disparity (DP) are based on the resulting predictions from GPT-3.5 and GPT-4 models.
@@ 8 Discussion and Conclusions
We introduce FWC, a novel coreset approach that generates synthetic representative samples along with sample-level weights for downstream learning tasks. FWC minimizes the Wasserstein distance between the distribution of the original datasets and that of the weighted synthetic samples while enforcing demographic parity. We demonstrate the effectiveness and scalability of FWC through experiments conducted on both synthetic and real datasets, as well as reducing biases in LLM predictions (GPT 3.5 and GPT 4). Future extensions include: (i) targeting different fairness metrics such as equalized odds (Hardt et al., 2016; Mishler et al., 2021), (ii) exploring privacy and explainability properties of FWC (Mohassel et al., 2020; Moshkovitz et al., 2020), (iii) exploring the theoretical fairness guarantees of the FWC by theoretically characterizing the fairness-utility tradeoff in non-i.i.d. settings  (Vidyasagar, 1997; Mohri & Rostamizadeh, 2008; Zhou et al., 2009), (iv) reformulating the optimization framework to target deep neural network pruning (Mussay et al., 2020; Ohib et al., 2022) and (v) investigate applications of fair synthetic data in the financial sector (Potluru et al., 2023).
Acknowledgements
ND is grateful to Alan Mishler for insightful discussions. Some of this work was performed while the first author was at JPMorgan Chase & Co. This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (“J.P. Morgan”), and is not a product of the Research Department of J.P. Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ Appendix A Details on Updating the Surrogate Function (line 4 of Algorithm 1)
To update the surrogate function, we need to solve problem (8), which is a huge-scale linear program with $O(n)$ constraints and $O(mn)$ nonnegative variables. In this work, we adapt FairWASP Xiong et al. (2024) for cases where $m\neq n$, as opposed to the scenario where $m=n$, which was tackled by Xiong et al. (2024). Before showing the main idea of the algorithm, we rephrase a useful lemma for doing linear minimization on $S_{n,m}\stackrel{{\scriptstyle\text{ def. }}}{{=}}\{P\in\mathbb{R}^{n\times m}:P\textbf{1}_{m}=\frac{1}{n}\cdot\mathbf{1}_{n},P\geq 0\}$.
@@
The proof of the above lemma is equivalent with that of Lemma 1 of Xiong et al. (2024) in the case when $m\neq n$, which can be extended directly. ∎
@@
in which $a_{j}^{\top}$ denotes the $j$-th row of $A$. Note that the problem (20) has much fewer decision variables than that of (19) and Lemma A.1 ensures the function $G(\cdot)$ is convex and has easily accessible function values and subgradients. Therefore, directly applying a cutting plane method has low per-iteration complexity and solves the problem (20) in linear time. We include the details on the cutting plane method in Section A.1 below. Finally, the primal optimal solution of (8) can be easily recovered from the dual optimal solution $\lambda^{\star}$ via solving $\max_{P\in S_{n,m}}\big{\langle}\sum_{j=1}^{h}\lambda_{j}^{\star}\mathbf{1}_{n}a_{j}^{\top}-C,P\big{\rangle}$, under the assumption that this problem has a unique minimizer, which almost always holds in practice for the computed $\lambda^{\star}$ and is also assumed by Xiong et al. (2024). In this way, we have shown how the problem (8) can be efficiently solved by applying a cutting plane method on its dual problem.
@@ A.1 Details of the Cutting Plane Method for Solving (20)
The cutting plane method is designed for convex problems where a separation oracle can be employed (Khachiyan, 1980). For any $\lambda\in\mathbb{R}^{m}$, a separation oracle operates by generating a vector $g$ which satisfies $g^{\top}\lambda\geq g^{\top}\lambda^{\star}$ for all $\lambda^{\star}$ in the set of optimal solutions. By repeatedly applying the separation oracle to cut down the potential possible feasible set, the cutting plane method progressively narrows down the feasible solution space until it reaches convergence. The specific steps of the cutting plane algorithm are detailed in Algorithm 2, with the key distinctions among different versions of this method lying in how lines 3 and 4 are implemented.
@@
For the problem (20), a separation oracle (line 4 in Algorithm 2) can directly use the vector of subgradients, which are efficiently accessible, as we mentioned in section 4. Given that we have shown (20) is a low-dimensional convex program with subgradient oracles, there exist many well-established algorithms that can be used. Suppose that the norm of an optimal $\lambda^{\star}$ is bounded by $R$, to the best of our knowledge, the cutting plane method with the best theoretical complexity is given by Jiang et al. (2020a), who proposed an improved cutting plane method that only needs $O((h\cdot\operatorname{SO}+h^{2})\cdot\log(hR/{\epsilon}))$ flops. Here SO denotes the complexity of the separation oracle. Note that here $h$ is at most $2|{\cal D}||{\cal Y}|$, which is far smaller than $n$ or $m$. Here we restate the Corollary 5 of Xiong et al. (2024) below for the overall time and space complexity of applying the cutting plane method in the case $m\neq n$.
Lemma A.2 (Essentially Corollary 5 of Xiong et al. (2024)).
@@
The above result is essentially Corollary 5 of Xiong et al. (2024) by slightly extending the proof to the general case $m\neq n$. Finally, in terms of the implementation, we follow Xiong et al. (2024) and use the analytic center cutting plane method.
@@ Appendix B Theoretical Proofs
This section includes the theoretical proofs for Section 2 and Section 5. We first show the Wasserstein distance upper bounds downstream disparity for general classifiers. We then show (i) the optimization problem can be reduced to optimizing over the $\widehat{X}$ rather than $\widehat{Z}$ (Lemma 3.1), (ii) the surrogate function is convex and a valid upper bound of the optimization objective (Lemma 5.2), (iii) our proposed algorithm converges to a first-order stationary point in $\widehat{X}$ (Theorem 5.3), and (iv) our proposed algorithm terminates in a finite amount of iterations (Theorem 5.4). We also prove the generalization bound for FWC performance in terms of Wasserstein distance and demographic parity to unseen datasets coming from the same (unknown) distribution the original dataset was sampled from. Finally, we include Section B.1 to note how the downstream learning using FWC can be broken down into two terms, which highlights the challenges in analyzing its theoretical properties.
@@
Proof of Proposition 2.1 The proof of the upper bound follows from the first part of the proof of the Kantorovich-Rubinstein duality (Santambrogio, 2015). In this work we follow the proof by Basso (2015) and Thickstun (2019), which consider the Lagrangian form of the 1-Wasserstein distance and express it in the following form:
$\mathcal{W}(p_{Z},p_{\widehat{Z}})=\sup_{f,g:f(x)+g(y)\leq\|x-y\|_{2}}\left|\mathbb{E}_{z\sim p_{Z}}f_{\theta}(z)-\mathbb{E}_{z\sim p_{\widehat{Z}}}f_{\theta}(z)\right|,$
$h(z)=\frac{f_{\theta}(z)}{L_{f_{\theta}}}.$
By definition, $h(z)$ is 1-Lipschitz. Again following Basso (2015) and Thickstun (2019), we know that for 1-Lipschitz functions the following holds:
$\left|\mathbb{E}_{z\sim p_{Z}}f_{\theta}(z)-\mathbb{E}_{z\sim p_{\widehat{Z}}}f_{\theta}(z)\right|$ $=L_{f_{\theta}}\left|\mathbb{E}_{z\sim p_{Z}}h(z)-\mathbb{E}_{z\sim p_{\widehat{Z}}}h(z)\right|$
$=L_{f_{\theta}}\left|\int_{\mathcal{Z}\times\mathcal{Z}}h(z_{1},z_{2})\pi(p_{Z},p_{\widehat{Z}})dz_{1},dz_{2}\right|$
@@
The results can be obtained by using the upper bound in Virmaux & Scaman (2018, Section 6.1), which shows that $L_{f_{\theta}}\leq L_{k}$ for $K$-layer MLPs with ReLu activations.
@@
Before proving it is an upper bound, we show some important properties of $F(C)$ as a function of $C$. Firstly, $F(C)$ is concave on $C$ because of the concavity of the minimum LP’s optimal objective on the objective vector. Secondly, since the feasible set of problem (8) is bounded, the optimal solution $F(C)$ is continuous with respect to $C$. Thirdly, due to the sensitivity analysis of LP (Bertsimas & Tsitsiklis, 1997), a supergradient of $F(C)$ at point $C$ is the corresponding optimal solution $P^{\star}$. Here the definition of supergradients for concave functions is analogous to the definition of subgradients for convex functions.
@@
Note that the first term is deterministic, as it is the result of the optimization problem (4). For the second term, we use the result from Fournier & Guillin (2015), which implies that with probability $1-\alpha$ and for the 1-Wasserstein distance:
@@
Note that the first term is deterministic, while the second one is not, as we need to account for the uncertainty of observing $n$ i.i.d. samples $\left\{Z_{i}\right\}_{i=1}^{n}$. For the second part, we use the Dvoretzky–Kiefer–Wolfowitz (DKW, Dvoretzky et al. 1956) inequality:

### v4

@@ 4.2 Updating Feature Vectors (Line 5)
$\min_{\widehat{X}_{j}\in{\cal X}:j\in[m]}\sum_{i\in[n]}\sum_{j\in[m]}c({Z}_{i},\widehat{Z}_{j})P_{ij}$ (12)
for $P=P_{k}^{\star}$, in which each component of $P$ is nonnegative and $\widehat{Z}_{j}=(\widehat{d}_{j},\widehat{X}_{j},\widehat{y}_{j})$, for the known fixed $\widehat{d}_{j}$ and $\widehat{y}_{j}$. Furthermore, the matrix $P$ is sparse, containing at most $n$ non-zeros (as when updating $P_{k}^{\star}$ for problem (8), see Appendix A). Moreover, problem (12) can be separated into the following $m$ subproblems,
$\min_{\widehat{X}_{j}\in{\cal X}}\sum_{i\in[n]}c(Z_{i},\widehat{Z}_{j})P_{ij}\ \text{, \ for }j\in[m]\ .$ (13)
@@ 6 An Alternative View: Generalized clustering algorithm
The minimizer $P^{\star}$ of (17) can be written in closed form. For each $i\in[n]$, let $C_{ij_{i}^{\star}}$ denote a smallest component on the $i$-th row of $C$. Then the components of a minimizer $P^{\star}$ can be written as $P_{ij}^{\star}=\frac{1}{n}\cdot\mathbf{I}(j=j_{i}^{\star})$ (where $\mathbf{I}$ is the indicator function). Hence, without fairness constraints, FWC corresponds to Lloyd’s algorithm for clustering. Specifically, Lloyd’s algorithm iteratively computes the centroid for each subset in the partition and subsequently re-partitions the input based on the closeness to these centroids [42]; these are the same operations FWC does in optimizing the surrogate function and solving problem (17). Thus, when $c(x,y)$ is correspondingly defined as $\|x-y\|_{1}$ or $\|x-y\|_{2}^{2}$, FWC corresponds to Lloyd’s algorithm applied to k-medians or k-means problems, except the centroids have fixed values for $\widehat{D}$ and $\widehat{Y}$ (see Section 3).
Comparison with k-means and k-medoids FWC and Lloyds’ algorithm for k-means or k-median share similar per-iteration complexity, with the main difference in complexity due to solving problem (8). We solve this problem efficiently by utilizing a variant of the FairWASP approach by [77] (Section 4.1), hence avoiding the usual complexity in solving optimal transport problems. As shown in Section 5, the leading term in the runtime complexity is $\mathcal{O}(nm)$, which comes from calculating and storing the cost matrix $C$. This level of complexity is the same as those in k-means and k-medoids. In addition, from our experiments we also see that the per-iteration complexity of FWC is roughly linear with the original dataset size $n$ (see the runtime experiment in Section 7 and Appendix C).
@@ 7 Experiments
We evaluate the performance of FWC  on 4 datasets widely used in the fairness literature [19]: (i) Adult [7], (ii) German Credit [28], (iii) Communities and Crime [64] and (iv) Drug [20]. For each dataset, we consider 3 different coreset sizes $m=5\%,10\%,20\%$ (apart from the Adult dataset, in which we select $m$ equal to 0.5%, 1% and 2% due to the large dataset size). We compare our approach with: (a) Fairlets and IndFair, two fair clustering approaches by [4] and [12], (b) K-Median Coresets, a coreset approach by [3], (c) k-means [42] and k-medoids [45, 58], two classic clustering approaches and (d) Uniform Subsampling of the original dataset. For FWC, we consider three different values of the fairness violation hyper-parameters $\epsilon$ for the optimization problem in (5). We compute the fairness-utility tradeoff by first training a 2-layer multilayer perceptron (MLP) classifier with ReLu activations on the coresets created by each approach and then evaluating the classifier demographic disparity (fairness) and AUC (utility). Figure 1 shows the model with the best fairness-utility tradeoff across the three coreset sizes $m$, for each approach. FWC  obtains equal or better fairness-utility tradeoffs (smaller disparity at the same level of utility, higher utility with the same disparity, or both) across all datasets, and performance remains competitive even when using a fairness pre-processing approach [34]. Appendix C.2 includes more experiments and details, which highlight that: (a) FWC  consistently achieves coresets that are closer in distribution to the original dataset with respect to the other methods and, although not natively minimizing clustering cost, also provide competitive performance for smaller datasets (Tables 3 and 4); (b) when added to the training data using the data augmentation schema proposed by [68, Section 2.1] FWC generally either increase the performance or reduce the demographic disparity in the downstream learning process.
@@
Figure 1: Top left: FWC runtime when changing the original dataset size $n$. Others: Fairness-utility tradeoff on real datasets for a downstream MLP classifier, selecting the model with the best fairness-utility tradeoff across three different coreset sizes $m$, with averages taken over 10 runs. FWC  consistently achieves a comparable/better tradeoff as shown by the Pareto frontier (dashed red line, computed over all models and coreset sizes), even when adjusting the other coresets with a fairness pre-processing technique [33]. See text and Appendix C.2 for more details.
Using FWC to improve fairness for LLM
@@
[76] evaluate GPT-3.5 and GPT-4 for fairness on predictive tasks for the UCI Adult dataset in a zero and few shot setting. We use a similar evaluation setup and use FWC in the few shot setting as examples and evaluate the results for the gender protected attribute. Specifically, we transform the tabular data into language descriptions, and ask GPT-3.5 Turbo and GPT-4 to perform classification tasks on it. We select $200$ samples to construct the test set and use a set of 16 samples found using FWC as examples. Further details on this experiment are provided in the appendix. The results are shown in Table 1. Examples provided by  FWC help reduce demographic disparity more than providing balanced few shot examples, while losing on predictive accuracy (note that the drop in accuracy is similar to the drop observed in [76] and is representative of the fairness-utility trade-off). Owing to the token limitation of LLM’s, these representative coresets can evidently be valuable to provide a small set of samples that can help mitigate bias. When accounting for the standard deviations for demographic parity in Table 1, FWC reduces the LLM bias when compared to zero shot prompting for GPT-4 across all runs. For GPT-3.5 Turbo, while the average disparity is reduced across runs, such consistency is indeed not observed, owing to a diverse set of outputs from the large language model. Due to limited availability of computational resources (associated with querying these models), we leave a more thorough evaluation across different datasets and models to future work.
@@
Table 1: Using the same setup as in [76], we use GPT-3.5 Turbo and GPT-4 LLM’s for fairness evaluations, with a test set of 200 samples with 0.5 base parity (${b_{p}}=0.5$). Few Shot - FWC is used to provide sixteen examples with weights to the model as examples. Accuracy and demographic disparity (DP) are based on the resulting predictions from GPT-3.5 and GPT-4 models.
@@ 8 Discussion and Conclusions
We introduce FWC, a novel coreset approach that generates synthetic representative samples along with sample-level weights for downstream learning tasks. FWC minimizes the Wasserstein distance between the distribution of the original datasets and that of the weighted synthetic samples while enforcing demographic parity. We demonstrate the effectiveness and scalability of FWC through experiments conducted on both synthetic and real datasets, as well as reducing biases in LLM predictions (GPT 3.5 and GPT 4). Future extensions include: (i) targeting different fairness metrics such as equalized odds [27, 49] as well as robustness of fairness-performance tradeoff over distribution shifts [48, 67], (ii) exploring privacy and explainability properties of FWC [50, 51], (iii) utilizing coresets for accelerating gradient descents algorithms and test their convergence [47, 70] (iv) reformulating the optimization framework to target deep neural network pruning [52, 54] and (v) investigate applications of fair synthetic data in the financial sector [61].
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JP Morgan”) and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
@@ Appendix A Details on Updating the Surrogate Function (line 4 of Algorithm 1)
To update the surrogate function, we need to solve problem (8), which is a huge-scale linear program with $O(n)$ constraints and $O(mn)$ nonnegative variables. In this work, we adapt FairWASP [77] for cases where $m\neq n$, as opposed to the scenario where $m=n$, which was tackled by [77]. Before showing the main idea of the algorithm, we rephrase a useful lemma for doing linear minimization on $S_{n,m}\stackrel{{\scriptstyle\text{ def. }}}{{=}}\{P\in\mathbb{R}^{n\times m}:P\textbf{1}_{m}=\frac{1}{n}\cdot\mathbf{1}_{n},P\geq\mathbf{0}\}$.
@@
The proof of the above lemma is equivalent with that of Lemma 1 of [77] in the case when $m\neq n$, which can be extended directly. ∎
@@
in which $a_{j}^{\top}$ denotes the $j$-th row of $A$. Note that the problem (20) has much fewer decision variables than that of (19) and Lemma A.1 ensures the function $G(\cdot)$ is convex and has easily accessible function values and subgradients. Therefore, directly applying a cutting plane method has low per-iteration complexity and solves the problem (20) in linear time. We include the details on the cutting plane method in Section A.1 below. Finally, the primal optimal solution of (8) can be easily recovered from the dual optimal solution $\lambda^{\star}$ via solving $\max_{P\in S_{n,m}}\big{\langle}\sum_{j=1}^{h}\lambda_{j}^{\star}\mathbf{1}_{n}a_{j}^{\top}-C,P\big{\rangle}$, under the assumption that this problem has a unique minimizer, which almost always holds in practice for the computed $\lambda^{\star}$ and is also assumed by [77]. In this way, we have shown how the problem (8) can be efficiently solved by applying a cutting plane method on its dual problem.
@@ A.1 Details of the Cutting Plane Method for Solving (20)
The cutting plane method is designed for convex problems where a separation oracle can be employed [36]. For any $\lambda\in\mathbb{R}^{m}$, a separation oracle operates by generating a vector $g$ which satisfies $g^{\top}\lambda\geq g^{\top}\lambda^{\star}$ for all $\lambda^{\star}$ in the set of optimal solutions. By repeatedly applying the separation oracle to cut down the potential possible feasible set, the cutting plane method progressively narrows down the feasible solution space until it reaches convergence. The specific steps of the cutting plane algorithm are detailed in Algorithm 2, with the key distinctions among different versions of this method lying in how lines 3 and 4 are implemented.
@@
For the problem (20), a separation oracle (line 4 in Algorithm 2) can directly use the vector of subgradients, which are efficiently accessible, as we mentioned in section 4. Given that we have shown (20) is a low-dimensional convex program with subgradient oracles, there exist many well-established algorithms that can be used. Suppose that the norm of an optimal $\lambda^{\star}$ is bounded by $R$, to the best of our knowledge, the cutting plane method with the best theoretical complexity is given by [31], who proposed an improved cutting plane method that only needs $O((h\cdot\operatorname{SO}+h^{2})\cdot\log(hR/{\epsilon}))$ flops. Here SO denotes the complexity of the separation oracle. Note that here $h$ is at most $2|{\cal D}||{\cal Y}|$, which is far smaller than $n$ or $m$. Here we restate the Corollary 5 of [77] below for the overall time and space complexity of applying the cutting plane method in the case $m\neq n$.
Lemma A.2 (Essentially Corollary 5 of [77]).
@@
The above result is essentially Corollary 5 of [77] by slightly extending the proof to the general case $m\neq n$. Finally, in terms of the implementation, we follow [77] and use the analytic center cutting plane method.
@@ Appendix B Theoretical Proofs
This section includes the theoretical proofs for Section 2 and Section 5. We first show the Wasserstein distance upper bounds downstream disparity for MLP networks (Proposition 2.1). We then show (i) the optimization problem can be reduced to optimizing over the $\widehat{X}$ rather than $\widehat{Z}$ (Lemma 3.1), (ii) the surrogate function is convex and a valid upper bound of the optimization objective (Lemma 5.2), (iii) our proposed algorithm converges to a first-order stationary point in $\widehat{X}$ (Theorem 5.3), and (iv) our proposed algorithm terminates in a finite amount of iterations (Theorem 5.4). We also prove the generalization bound for FWC performance in terms of Wasserstein distance and demographic parity to unseen datasets coming from the same (unknown) distribution the original dataset was sampled from. Finally, we include Section B.1 to note how the downstream learning using FWC can be broken down into two terms, which highlights the challenges in analyzing its theoretical properties.
Lemma B.1.
Let $Z=(X,Y),\widehat{Z}=(\widehat{X},\widehat{Y})\in{\cal Z}=({\cal X}\times{\cal Y})$ be two pairs of random variables with joint distributions $p_{Z}$ and $p_{\widehat{Z}}$ and marginal distributions $p_{X}$, $p_{Y}$ and $p_{\widehat{X}}$ and $p_{\widehat{Y}}$ respectively. Let $\Pi(Z,\widehat{Z})$ indicate the set of all joint probability distributions over the product space ${\cal Z}\times{\cal Z}$ that admit marginal and conditional distributions over ${\cal X}$ and ${\cal Y}$. For any non-negative cost operator $c$:
$\mathcal{W}_{c}(p_{X},p_{\widehat{X}})\leq\mathcal{W}_{c}(p_{Z},p_{\widehat{Z}}).$
@@
Proof of Lemma B.1 Let $\Pi_{X}(X,\widehat{X})$ indicate the set of all marginal (joint) probability distributions over the product space ${\cal X}\times{\cal X}$. For any $\pi\in\Pi(Z,\widehat{Z})$ and the corresponding $\pi_{x}\in\Pi_{X}(X,\widehat{X})$ we have that:
@@
$\geq\min_{\pi_{x}\in\Pi_{x}(x,\widehat{x})}c(x,\widehat{x})d\pi_{x}(x,\widehat{x})=\mathcal{W}_{c}(p_{X},p_{\widehat{X}})$
As this is valid for any $\pi\in\Pi(Z,\widehat{Z})$, select:
[…]
Then the left-hand-side of (22) is equal to $\mathcal{W}_{c}(p_{Z},p_{\widehat{Z}})$, hence proving that $\mathcal{W}_{c}(p_{Z},p_{\widehat{Z}})\geq\mathcal{W}_{c}(p_{X},p_{\widehat{X}})$.
Proof.
Proof of Proposition 2.1 The proof of the upper bound follows from the first part of the proof of the Kantorovich-Rubinstein duality [66]. In this work we follow the proof by [6, 71], which consider the Lagrangian form of the 1-Wasserstein distance and express it in the following form:
$\mathcal{W}(p_{(X,D)},p_{(\widehat{X},\widehat{D})})=\sup_{f,g:f(x)+g(y)\leq\|x-y\|_{2}}\left|\mathbb{E}_{(x,d)\sim p_{(X,D)}}f_{\theta}(x,d)-\mathbb{E}_{(x,d)\sim p_{(\widehat{X},\widehat{D})}}f_{\theta}(x,d)\right|,$
$h(x,d)=\frac{f_{\theta}(x,d)}{L_{f_{\theta}}}.$
By definition, $h(x,d)$ is 1-Lipschitz. Again following [6, 71], we know that for 1-Lipschitz functions the following holds:
$\left|\mathbb{E}_{(x,d)\sim p_{(X,D)}}f_{\theta}(x,d)-\mathbb{E}_{(x,d)\sim p_{(\widehat{X},\widehat{D})}}f_{\theta}(x,d)\right|=L_{f_{\theta}}\left|\mathbb{E}_{(x,d)\sim p_{(X,D)}}h(x,d)-\mathbb{E}_{(x,d)\sim p_{(\widehat{X},\widehat{D})}}h(x,d)\right|$
[…]
$\quad\leq L_{f_{\theta}}\mathcal{W}_{1}(p_{(X,D)},p_{(\widehat{X},\widehat{D})})\leq L_{f_{\theta}}\mathcal{W}_{1}(p_{Z},p_{\widehat{Z}}),$
where the last inequality is due to Lemma B.1. The result can be obtained by using the upper bound in [75, Section 6.1], which shows that $L_{f_{\theta}}\leq L_{k}$ for $K$-layer MLPs with ReLu activations.
@@
Before proving it is an upper bound, we show some important properties of $F(C)$ as a function of $C$. Firstly, $F(C)$ is concave on $C$ because of the concavity of the minimum LP’s optimal objective on the objective vector. Secondly, since the feasible set of problem (8) is bounded, the optimal solution $F(C)$ is continuous with respect to $C$. Thirdly, due to the sensitivity analysis of LP [8], a supergradient of $F(C)$ at point $C$ is the corresponding optimal solution $P^{\star}$. Here the definition of supergradients for concave functions is analogous to the definition of subgradients for convex functions.
@@
Note that the first term is deterministic, as it is the result of the optimization problem (4). For the second term, we use the result from [22], which implies that with probability $1-\alpha$ and for the 1-Wasserstein distance:
@@
Note that the first term is deterministic, while the second one is not, as we need to account for the uncertainty of observing $n$ i.i.d. samples $\left\{Z_{i}\right\}_{i=1}^{n}$. For the second part, we use the Dvoretzky–Kiefer–Wolfowitz (DKW, [17]) inequality: